OpenStudy (tayidaniels):

Evaluate −6 − (−3). 18 −9 3 −3

OpenStudy (joker):

-6 - (-3) = -6 + 3 What is -6 + 3 = ?

OpenStudy (tayidaniels):

@Joker -3

OpenStudy (kgrendel0324):

exactly

sammixboo (sammixboo):

@kgrendel0324 Welcome to OpenStudy :) I want to let you know that direct answers aren't allowed on OpenStudy. We love to guide users to the answer, and teach them.

OpenStudy (tayidaniels):

@kgrendel0324 Thank you both for the help!